Shift report 20011013 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 13 2001 - 23:57:07 EDT

  • Next message: Apache: "Shift report 20011014 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Oct 13 23:57:07 2001, Shift: 16:00-24:00
    Supervisor: CEJ, with: DO
    Log book#9 pages 148 - 149
    
    Shift Summary:
    ------------------------------------------------------------
    1. Short store in the beginning of the run. Problems
       steering, low ZDC rates at BRAHMS and STAR. Dumped
       beam at 17.30. Cryo problem, problem in AGS to RHIC
       transfer line, changing of a foil in the tandem,
       quench link interlock. Several attempts to store and 
       ramp beam...no luck this evening. Last night the 
       intensity looked like the alpes - this evening it was 
       more like sjaelland (very flat!).
    
    2. Two short data runs:
       5313 (T3 anode3 at 900V) & 5314 
       Pedestal run 5315, TPC pedestal run 5316
    
    3. YKL tried to fix tile 33, we need data to see the 
       result.	
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 13 2001 - 23:57:51 EDT